OpenStudy (anonymous):

The following table shows the probability distribution for a discrete random variable. X- 12,15,17,20,22,24,25,30 P(X)- 0.07,0.21,0.17,.025,0.05,0.04,0.13,0.08 The mean of the discrete random variable X is 19.59. What is the variance of X? Round your answer to the nearest hundredth.

OpenStudy (amistre64):

n*p*q ?? or is that for a binomial

OpenStudy (anonymous):

I have no idea that is all the pieces of problem i got,..

OpenStudy (amistre64):

\[var(X)=\sum_{n}{(n-E(x))^2*p_X(n)}\] ..... whatever that means

OpenStudy (amistre64):

please do not bypass the filters in order to use profanity. We try to keep this a family freindly place to interact in.

OpenStudy (amistre64):

http://www.statisticslectures.com/topics/variancestandarddeviationdiscrete/ this looks more readable to me

OpenStudy (amistre64):

lets use this setup to see what we can come up with: x P(x) x^2 x^2*P(x) 12 0.07 144 15 0.21 225 17 0.17 289 20 .025 400 22 0.05 484 24 0.04 576 25 0.13 625 30 0.08 900 is that P(20) correct?

OpenStudy (anonymous):

that looks good

OpenStudy (amistre64):

0.07* 144 +0.21* 225 +0.17* 289 + .025* 400 + 0.05* 484+ 0.04* 576+ 0.13* 625+ 0.08* 900 = 316.95 if i assume P(20) is a typo i get 0.07* 144 +0.21* 225 +0.17* 289 + 0.25* 400 + 0.05* 484+ 0.04* 576+ 0.13* 625+ 0.08* 900 = 406.95 but im not sure if we are done lol

OpenStudy (amistre64):

looks like we need to subtract the square of the E(x) from that as well

OpenStudy (amistre64):

19.56^2 = 382.5936 so subtract that from the correct version of P(20) and it should be good

OpenStudy (anonymous):

yea. it was 0.25

OpenStudy (amistre64):

whew!! good :) the process seems a bit long and cumbersome; but a calculator should help 406.95 - 382.5936 ----------- var(x)

OpenStudy (anonymous):

24.356

OpenStudy (amistre64):

with any luck, yes

OpenStudy (anonymous):

outstanding.

OpenStudy (amistre64):

dbl chk my math to be sure ;)